an helicopter hoovers clear on the ground issuing slip diameter 20 meters at speed of 20 m/s. Determine the mass flow rate of the air passing through the rotor and the mass of the helicopter?
1
Expert's answer
2020-02-21T09:59:37-0500

As per the question,

It is given that the diameter of the slip =20m

So, the radius of the slip =diameter/2=10m

Speed =20m/sec

The density of the air"=1.225 kg\/m^3"

So, mass flow rate = area covered by the slip x velocity of the slip

"=\pi r^2\times v=\dfrac{22\times 10\times 10\times 20}{7}=6.285\times10^3m^3/sec"

We know that mass of the helicopter= density of the air x flow rate

="6.285\times1.225\times 10^3=7.699\times10^3 kg=7699kg"